OpenStudy (anonymous):

whats 4|-10| - 3|-2| explain how you got the answer please

OpenStudy (anonymous):

Okay, first you figure out the absolute values. |-10| is the same thing as 10. That goes for |-2| as well, which is 2. You plug those values back into the equation and you get the following 4(10)-3(2)=? 4(10)=40 and 3(2) =6 Plug those in again, and you get 40-6=34

OpenStudy (anonymous):

Wait wait, that isn't the complete answer. You actually have 2 answers. I'll show that now.

myininaya (myininaya):

2 answers?

OpenStudy (anonymous):

Well, more like 4. Sorry, trying to be helpful. When you have an absolute value equation, such as |x|, the x equals both -x and +x. Since your equation is 4|-10|-3|2| you actually have 4 possible answers. 1. 4(10)-3(2)=34 2. 4(-10)-3(-2)=-34 3. 4(10)-3(-2)=46 4. 4(-10)-3(2)=-46 Your answers are +- 46 and +-34. That translates into plus or minus 46 and plus or minus 34

OpenStudy (anonymous):

\[\pm46 and \pm34\]

myininaya (myininaya):

34 is the only answer. if you have |x|=2 and you want to solve this for x, then x=-2 and x=2

OpenStudy (anonymous):

Myininaya is right, I was overthinking that :P